OpenStudy (anonymous):

What is the arc measure of an arc with length 4.189 cm and radius equal to 3 cm. A. 480 B. 53 C. 82 D. 80

Directrix (directrix):

D. 80

Directrix (directrix):

4.189/[2 π 3] = x/360 6 π x = 360 *4.189 x = 80.0 degrees approx I think you wanted the central angle measure of the angle that subtends that arc.

OpenStudy (anonymous):

thanks

Directrix (directrix):

Any time.